y=2x^2-9x-5 how do I find the 0 of the function.
Thanks in advance...

Answer by waynest(281)

factor:
2x^2 - 9x - 5 = y
(2x + 1)(x - 5)
check:
2x^2 - 10x + x - 5
2x^2 - 9x - 5
zeros:
2x + 1 = 0
2x = -1
2x/2 = -1/2
x = -1/2
x - 5 = 0
x - 5 + 5 = 5
x = 5